OpenStudy (anonymous):

Factor (5x^2-16x+11)

OpenStudy (anonymous):

What is the quickest way to go about factoring this? The 5 is throwing me off.

OpenStudy (zzr0ck3r):

5x^2 -5x-11x+11 = 5x(x-1)-11(x-1) = (5x-11)(x-1)

OpenStudy (zzr0ck3r):

note i used the AC method 5*11 = 55 = -11*-5 and -11+-5 = -16

OpenStudy (anonymous):

Given a polynomial with degree 2, we need to find numbers m, n such that, for some P(x): \[ P(x)=ax^2+bx+c\\ mn=ac\\ m+n=b \]So, we take these, and do the following: \[ P(x)=ax^2+mx+nx+c \]And factor. Like @zzr0ck3r 's answer.

jimthompson5910 (jim_thompson5910):

Optionally you can set the expression equal to zero and then use the quadratic formula to get the two solutions x = 1 or x = 11/5 Now rearrange so that 0 is on the right side and then use the zero product property x = 1 or x = 11/5 x = 1 or 5x = 11 x - 1 = 0 or 5x - 11 = 0 (x - 1)(5x - 11) = 0 So the expression factors to (x - 1)(5x - 11)

OpenStudy (zzr0ck3r):

just a note, we do not have to set it equal to anything.

OpenStudy (zzr0ck3r):

right?

jimthompson5910 (jim_thompson5910):

True, but I find this method straight to the point (and you avoid guess and check)

OpenStudy (zzr0ck3r):

I c

jimthompson5910 (jim_thompson5910):

Either way works

OpenStudy (anonymous):

Thanks for the help guys. Wish I could give more medals.

OpenStudy (zzr0ck3r):

Yeah, was more just checking that I was ok doing what I did, but light bulb....

OpenStudy (zzr0ck3r):

jim does not need any more:p and I gave wolf one:)

jimthompson5910 (jim_thompson5910):

lol idc about them so don't worry

OpenStudy (anonymous):

Oy @jim_thompson5910 remember that you also have to multiply it by the GCD of all of the coefficients.

jimthompson5910 (jim_thompson5910):

in this case, the GCD is 1, so it's still the same answer